How do you graph (x+4)²+(y-2)²=25? What is the area of the enclosed region?
Found 2 solutions by Edwin McCravy, nyc_function:
Answer by Edwin McCravy(20078)

(x+4)² + (y-2)² = 25

is in the form:

(x-h)² + (y-k)² = r²

which is the equation of a circle with
center (h,k) = (-4,2) and radius 5.



Its area is given by the formula A = pr² = p(5)² = p(25) = 25p   


Edwin



Answer by nyc_function(2741)

This is a circle centered at (-4,2) and
has a radius of 5. Plot the point (-4,2)
and move a distance of 5 units from
the point up, down, left and right.
Then connect all the points forming
a circle.
Since the radius is 5, use
A = pi(radius)^2 to find area.
A = pi(5)^2
A = 25pi